OpenStudy (anonymous):

a camera manufacturer spends $1,800 each day for overhead expenses plus $9 per camera for labor and materials. the cameras sell for $15 each. a) how many cameras must the company sell in one day to equal its daily costs ? b) if the manufacturer can increase production by 50 cameras per day, what would their daily profit be?

OpenStudy (anonymous):

1800+9+15=1824 24x=1824 A x=76 cameras B 50x24=1200+1800=$3000

OpenStudy (anonymous):

how did u get the 24 ?

OpenStudy (anonymous):

sorry 9 and 15

OpenStudy (anonymous):

oh ok thanks! (:

OpenStudy (anonymous):

wait 1800 24x=1800 A x=75 cameras

OpenStudy (anonymous):

B is 75+50=125x24=3000+1800=$4800